A 5 days remaining in your free trial. Purchaseaccessnow An instructor who taught two sections of engineering statistics last term, the first with 25 students and the second with 40, decided to assign a term project. After all projects had been turned in, the instructor randomly ordered them before grading. Consider the first 15 graded projects (a) What is the probability that exactly 10 of these are from the second section? (Round your answer to four decimal places.) (b) What is the probability that at least 10 of these are from the second section? (Round your answer to four decimal places.) (e) What is the probability that at least 10 of these are from the same section? (Round your answer to four decimal places.) (d) What are the mean value and standard deviation of the number among t section? (Round your mea mean standard deviation hese 15 that are from the second n to the nearest whole number and your standard deviation to three decimal places.) projects projects (e) What are the mean value and standard deviation of the number of projects not among these first 15 that are from the second section? (Round your mean to the nearest whole number and your standard deviation to three decimal places.) mean standard deviation projects projects
Answer #1

a)

from hypergeometric distribution: probability that exactly 10 from second edition P(X=10)=\binom{25}{5}\binom{40}{10}/\binom{65}{15}=0.2172

b)

P(at least 10 from second edition)=P(X>=10)=0.4406

c)

P(at least 10 from second edition)=P(at least 10 from 1st )+P(at least 10 from 2nd edition )=0.4532

d)

mean =nk/N =15*40/65=9

std deviation =(nk/N*(1-k/N)*(N-n)/(N-1))1/2 =1.665

e)

mean=15-9=6

std deviation =1.665

( please revert fr any clarification required)